OpenStudy (anonymous):

The expression log (x^n/ sqredroot z) is equivalent to ... muiltiple choice just want to know how to do it :] thanks

myininaya (myininaya):

do you want to expand log(x/y)=log(x)-logy see if you can use this and then we will talksome more

OpenStudy (anonymous):

Thats it i dont know how to even do that the muiltiple choice answers are set up as A) log (nx) - log (1/2z) B) Log (nx) - log(sz) and etc

myininaya (myininaya):

can you try to use log(x/y)=log(x)-log(y)

OpenStudy (anonymous):

when i set it up is it log x^n/\[\sqrt{z}\]=log x^n)-log(\[\sqrt{z}\] ...then how do i go from there if that is correct.